# Towards Underwater Camouflaged Object Tracking: Benchmark and Baselines

### v2

@@ Abstract
Over the past decade, significant progress has been made in visual object tracking, largely due to the availability of large-scale datasets. However, existing tracking datasets are primarily focused on open-air scenarios, which greatly limits the development of object tracking in underwater environments. To bridge this gap, we take a step forward by proposing the first large-scale multimodal underwater camouflaged object tracking dataset, namely UW-COT220. Based on the proposed dataset, this paper first comprehensively evaluates current advanced visual object tracking methods and SAM- and SAM2-based trackers in challenging underwater environments. Our findings highlight the improvements of SAM2 over SAM, demonstrating its enhanced ability to handle the complexities of underwater camouflaged objects. Furthermore, we propose a novel vision-language tracking framework called VL-SAM2, based on the video foundation model SAM2. Experimental results demonstrate that our VL-SAM2 achieves state-of-the-art performance on the UW-COT220 dataset. The dataset and codes can be accessible at https://github.com/983632847/Awesome-Multimodal-Object-Tracking.
@@ 1 Introduction
Visual object tracking (VOT) involves continuously locating a target object within a video sequence and has applications in autonomous vehicles xie2024diffusiontrack ; wang2024elysium ; wang2024event , surveillance zhang2024awesome ; cai2024hiptrack , and robotics zhang2022webuav ; zhang2023all . Its significance lies in enabling machines to perceive and interpret dynamic environments, supporting tasks like motion analysis and decision-making li2024visual ; hu2024multi ; peng2024vasttrack . Although significant progress has been made in terrestrial and open-air scenarios huang2019got ; fan2021lasot , tracking in underwater environments remains challenging due to factors like visual camouflage, light scattering, and low contrast, which limit the effectiveness of conventional algorithms zhang2024webuot ; alawode2023improving ; alawode2022utb180 . Consequently, there is a pressing need for specialized datasets and methods to tackle the complexities of underwater tracking, especially when objects blend with their surroundings, known as camouflaged objects. However, despite its importance, underwater camouflaged object tracking remains an unexplored field.
Current VOT methods can be broadly categorized into several types. Traditional correlation filter-based methods zhang2019robust ; ge2020cascaded ; ge2019distilling ; zhang2020accurate utilize efficient convolution operations to track objects, whereas Siamese-based methods bertinetto2016fully ; li2019siamrpn++ ; li2018high employ dual-stream networks to learn a similarity function between the target object and the search region. Transformer-based methods, such as OSTrack ye2022joint , SeqTrack chen2023seqtrack , and ARTrack wei2023autoregressive , leverage the attention mechanism to model complex dependencies in sequences, resulting in enhanced tracking precision. In recent years, Mamba-based approaches like Mamba-FETrack huang2024mamba , MambaVT lai2024mambavt , and MambaTrack zhang2024mambatrack have emerged, offering robust performance through feature enhancement and advanced attention mechanisms. In addition to these methods, SAM-based models zhang2023comprehensive ; zhang2024segment have also gained traction. These models, such as SAM-DA fu2023sam , Tracking Anything yang2023track , SAM-Track cheng2023segment , and the latest SAMURAI-L yang2024samurai , focus on the precise segmentation of objects across frames, making them particularly suitable for challenging environments like underwater scenarios where object boundaries are difficult to discern.
Figure 1: Some examples come from our UW-COT220 dataset. We provide bounding box and language description annotations for a camouflaged underwater target object in each video sequence.
Table 1: Comparison of the constructed UW-COT220 with existing camouflaged object tracking dataset (i.e., COTD guo2024cotd ) and video camouflaged object detection datasets (i.e., CAD bideau2016s and MoCA-Mask cheng2022implicit ). UW-COT220 is the first multimodal underwater camouflaged object tracking dataset.
@@
Annotation Link
@@
CAD bideau2016s 2016 9 6 30 93 218 836 191 Mask URL
MoCA-Mask cheng2022implicit 2022 87 45 23 264 1,296 23 K 23 K Mask URL
COTD guo2024cotd 2024 200 20 - 400 - 80 K 80 K BBox URL
UW-COT220 (Ours) 2024 220 96 10 722 7,448 159 K 159 K BBox+Mask URL
Following the latest large-scale underwater object tracking dataset (i.e., WebUOT-1M zhang2024webuot ), we take a step forward and construct the first multimodal underwater camouflaged object tracking dataset, UW-COT220 (see Fig. 1 and Tab. 1), as a benchmark for evaluating and developing advanced underwater camouflaged object tracking methods. The dataset consists of 220 underwater video sequences, spanning 96 categories, with approximately 159,000 frames. We annotate each video with a language description to promote the exploration of multimodal underwater camouflaged object tracking. As a single-object visual tracking task, we provide bounding box annotations for the camouflaged underwater objects in each frame. Additionally, using manually annotated bounding boxes as prompts, we generate pseudo mask annotations for the camouflaged underwater objects by leveraging advanced interactive segmentation models (e.g., HQ-SAM ke2024segment and SAM2 ravi2024sam ). We conduct pioneering experiments on the constructed dataset to evaluate the performance of state-of-the-art visual object tracking methods and the latest advancements in image and video segmentation. By releasing UW-COT220, we aim to inspire interest in exploring various underwater vision tasks and to foster the development of advanced methods for underwater camouflaged object tracking.
@@ 2 UW-COT220 Dataset
Our goal is to construct a large-scale underwater camouflaged object tracking dataset that involves a rich variety of categories and various real underwater scenes for evaluating and developing general underwater camouflaged object tracking methods. To achieve this, we collect underwater videos from video-sharing platforms (e.g., YouTube [^1] https://www.youtube.com/) and existing tracking datasets (e.g., WebUOT-1M zhang2024webuot and VastTrack peng2024vasttrack ), and we filter out 220 videos that contain underwater camouflaged objects. Following existing datasets huang2019got ; fan2019lasot ; zhang2022webuav ; peng2024vasttrack ; zhang2024webuot , we provide a bounding box annotation for the camouflaged object in each frame, represented as $[x,y,w,h]$, where $x$ and $y$ denote the coordinates of the top-left corner of the object, and $w$ and $h$ indicate the width and height of the object. In each video, we annotate an English sentence describing the color, behavior, attributes, and surroundings of the target. Furthermore, to enhance the accuracy of object representation, we use segmentation foundation models (e.g., HQ-SAM ke2024segment and SAM2 ravi2024sam ) to generate pseudo mask annotations for the camouflaged objects. Some examples and detailed statistics of the UW-COT220 dataset are provided in Fig. 1 and Tab. 1. To the best of our knowledge, UW-COT220 is the first large-scale benchmark for underwater camouflaged object tracking, featuring a diverse set of 96 categories and various challenging underwater scenes.
@@ 3 Proposed Method
In this section, we introduce VL-SAM2, a new vision-language tracking framework for multimodal underwater camouflaged object tracking, as shown in Fig. 2. The proposed VL-SAM2 is based on the latest video foundation model, SAM2, and extends its capabilities to challenging underwater scenarios. VL-SAM2 consists of three main components: a visual branch (i.e., SAM2), a language branch, and a training-free motion-aware target prediction (MATP) zhang2024webuot module.
@@
The input to VL-SAM2 consists of video frames and a language description. First, we use an image encoder and a language encoder to extract features from both modalities. In the first frame, the initial bounding box (BBox) is encoded by the prompt encoder to obtain the visual embedding. The language embedding is processed by a linear projection to match the dimension of the visual embedding. Then, we concatenate the visual and language embeddings as sparse embeddings and feed them into the mask decoder. In subsequent frames, we adopt a video propagation manner similar to SAM2 to obtain the prediction result for each frame. The MATP, based on Kalman filtering kalman1960new , is used to alleviate the model drift caused by similar distractors around the true target. This module was proposed in zhang2024webuot . Inspired by yang2024samurai , we further apply MATP to update the memory of SAM2. We strongly recommend readers refer to zhang2024webuot and yang2024samurai for more details.
@@
We use Hiera-L ryali2023hiera as the image encoder and CLIP ViT-B/32 radford2021learning as the language encoder. During training, we freeze the parameters of both the image encoder and the language encoder, and fine-tune the other modules. This allows us to train the proposed VL-SAM2 on an Ubuntu server with 8 A6000 GPUs. The current version has been trained only on the Ref-Youtube-VOS seo2020urvos dataset and then evaluated in a zero-shot manner on the UW-COT220 dataset.
@@
Figure 3: Overall performance on the UW-COT220 dataset using AUC, Pre, cAUC, and nPre scores.
@@
Table 2: Impact of using different points (within the initial target box) as the point prompt for SAM2 ravi2024sam . We report AUC, nPre, Pre, cAUC, and mACC scores on the UW-COT220 dataset.
  Method Point type AUC (%) Pre (%) nPre (%) cAUC (%) mACC (%)
SAM2-tiny Center point 53.1 53.7 60.1 52.2 53.4
Random point 37.7 37.2 41.6 36.4 37.3
SAM2-small Center point 52.8 53.6 60.2 52.0 53.1
Random point 36.3 35.3 41.0 35.2 36.0
SAM2-base plus Center point 55.0 56.0 62.2 54.2 55.4
@@
Table 3: Comparison of different models for SAM2 ravi2024sam on the UW-COT220 dataset.
  Method Size (M) Speed (FPS) AUC (%) Pre (%) nPre (%) cAUC (%) mACC (%)
SAM2-tiny 38.9 47.2 53.1 53.7 60.1 52.2 53.4
SAM2-small 46.0 43.3 52.8 53.6 60.2 52.0 53.1
SAM2-base plus 80.8 34.8 55.0 56.0 62.2 54.2 55.4
SAM2-large 224.4 24.2 60.6 63.5 67.3 60.0 61.3
4 Experimental Results
Main Results. We evaluate SAM-based tracking methods (i.e., SAM-DA fu2023sam , and Tracking Anything yang2023track ), SAM2, and several current SOTA VOT methods (i.e., OSTrack ye2022joint , SeqTrack chen2023seqtrack , ARTrack wei2023autoregressive , and SAMURAI-L yang2024samurai ) on the proposed UW-COT220 dataset. The results are shown in Fig. 3. Our observations are as follows: 1) Our VL-SAM2 achieves the best performance, surpassing the SOTA trackers, including SAMURAI-L yang2024samurai and SAM2. 2) The top three trackers (VL-SAM2, SAMURAI-L, and SAM2) are all SAM2-based, highlighting the promising trend of using foundation models for tracking. 3) SAM2-based methods outperform SAM-based trackers (SAM-DA and Tracking-Anything) on UW-COT220, which can be attributed to a series of improvements SAM2 introduces over SAM for video and image tasks, such as improving temporal consistency, robustness to occlusions, feature embeddings, computational efficiency, motion estimation accuracy, generalization to new domains, and integration of contextual information.
Impact of Prompt. We take SAM2 as our baseline tracker to explore the impact of different ways of point prompts (i.e., center point and random point within the initial target box). From Tab. 2, we find that using the center point as a prompt yields significantly better results than using random points. This implies that for the interactive segmentation model SAM2, the quality of the input prompt plays a crucial role in determining its performance. This also motivates us to explore the use of more flexible prompts, e.g., language descriptions, in the context of underwater camouflaged object tracking.
Impact of Model Size. The results in Tab. 3 demonstrate that larger model sizes generally lead to better performance, but the speed of the model decreases significantly. We also discovered an interesting phenomenon: when the number of model parameters is relatively small, a smaller model (e.g., SAM2-tiny) can even outperform a larger model (e.g., SAM2-small). We suspect this may be due to overfitting, or that small models (e.g., SAM2-small) are more sensitive to the quality of the training data ravi2024sam .
Impact of Language Encoder. In Tab. 4, we explore the impact of using different language encoders in the proposed VL-SAM2 framework. T5 (Text-to-Text Transfer Transformer) raffel2020exploring is a large-scale language model that models all natural language processing tasks as text-to-text problems, while CLIP radford2021learning is a contrastive language-image pre-training model. The results demonstrate that using CLIP as the language encoder outperforms T5. We hypothesize that this is due to CLIP’s text encoder being aligned with the visual encoder in the feature space over a vast amount of data, making it more suitable for the VOT task.
@@ 4 Experimental Results
  Method Language Encoder AUC (%) Pre (%) nPre (%) cAUC (%) mACC (%)
VL-SAM2 T5 raffel2020exploring 54.6 54.8 59.3 53.9 55.1
VL-SAM2 CLIP radford2021learning 71.3 75.2 78.2 70.8 72.4

### v3

# Underwater Camouflaged Object Tracking Meets Vision-Language SAM2

@@ Abstract
Over the past decade, significant progress has been made in visual object tracking, largely due to the availability of large-scale datasets. However, these datasets have primarily focused on open-air scenarios and have largely overlooked underwater animal tracking—especially the complex challenges posed by camouflaged marine animals. To bridge this gap, we take a step forward by proposing the first large-scale multi-modal underwater camouflaged object tracking dataset, namely UW-COT220. Based on the proposed dataset, this work first comprehensively evaluates current advanced visual object tracking methods, including SAM- and SAM2-based trackers, in challenging underwater environments, e.g., coral reefs. Our findings highlight the improvements of SAM2 over SAM, demonstrating its enhanced ability to handle the complexities of underwater camouflaged objects. Furthermore, we propose a novel vision-language tracking framework called VL-SAM2, based on the video foundation model SAM2. Experimental results demonstrate that our VL-SAM2 achieves state-of-the-art performance on the UW-COT220 dataset. The dataset and codes are available at here.
@@ 1 Introduction
Visual object tracking (VOT) involves continuously locating a target object within a video sequence and has applications in animal monitoring and conservation [34, 36], autonomous vehicles [31, 25], and surveillance [39, 8], etc. Its significance lies in enabling machines to perceive and interpret dynamic environments, supporting tasks like animal behavioral analysis [38] and decision-making [22]. Although substantial progress has been achieved in terrestrial and open-air scenarios [16, 9], tracking in underwater environments remains highly challenging, largely limiting the effectiveness of conventional algorithms [38, 2, 1]. Consequently, there is a pressing need for specialized datasets and methods to tackle the complexities of underwater tracking, especially when objects blend with their surroundings, known as camouflaged objects [4, 6]. However, despite its importance, underwater camouflaged object tracking remains an unexplored field.
Underwater object tracking presents numerous unique challenges due to light attenuation [38, 2], scattering effects [1, 22], and frequent camouflage scenarios [6, 4], which significantly degrade the performance of conventional tracking paradigms. Correlation filter-based methods [32, 12, 11, 33], while computationally efficient, fail to handle the dynamic appearance variations caused by underwater optical distortions. Siamese networks [3, 20] struggle with feature representation learning in low-contrast and turbid environments. Transformer-based approaches [31, 5, 28], though powerful in modeling long-range dependencies, often suffer from attention distraction due to cluttered underwater backgrounds. Even the emerging Mamba-based methods [15, 19, 40], despite their efficient sequence modeling, lack the specialized capability to address underwater-specific degradations. In contrast, SAM- and SAM2-based trackers [35, 37, 10, 30, 29] demonstrate superior robustness by leveraging their pre-trained visual foundation model’s strong generalization ability and prompt-guided segmentation mechanism. This enables effective handling of camouflaged marine animals through semantic-aware feature extraction and adaptive multi-scale processing, making SAM- and SAM2-based approaches particularly suitable for challenging underwater tracking scenarios.
Figure 1: A glance at some examples covering diverse underwater scenes in our UW-COT220 dataset. We provide extensive annotations for underwater camouflaged target objects in video sequences, including bounding boxes, masks, and language descriptions.
Table 1: Comparison of the constructed UW-COT220 with existing camouflaged object tracking dataset (i.e., COTD [14]) and video camouflaged object detection/segmentation datasets (i.e., CAD [4] and MoCA-Mask [6]). UW-COT220 is the first multi-modal underwater camouflaged object tracking dataset.
@@
Visual
annotation
Absent
label
@@
Open
source
CAD [4] 2016 9 6 30 93 218 836 191 Mask
MoCA-Mask [6] 2022 87 45 23 264 1,296 23 K 23 K Mask
COTD [14] 2024 200 20 - 400 - 80 K 80 K BBox
UW-COT220 (Ours) 2025 220 96 10 722 7,448 159 K 159 K BBox+Mask
Recently, several video camouflaged object detection and segmentation datasets have been proposed [4, 6, 14], but they primarily focus on open-air scenarios. Following the latest large-scale underwater object tracking dataset, i.e., WebUOT-1M [38], we take a step forward and construct the first multi-modal underwater camouflaged object tracking dataset, UW-COT220 (see Fig. 1 and Tab. 1), as a benchmark for evaluating and developing advanced underwater camouflaged object tracking methods. The dataset consists of 220 underwater video sequences, spanning 96 categories, with approximately 159 K frames. We annotate each video with a language description to promote the exploration of multi-modal underwater camouflaged object tracking. As a single-object visual tracking task, we provide bounding box annotations for the camouflaged underwater objects in each frame. Additionally, using manually annotated bounding boxes as prompts, we generate mask annotations for the camouflaged underwater objects. To advance research in the related field, we propose a simple yet effective VL tracking framework, called VL-SAM2. By releasing our dataset and codes, we aim to inspire the exploration of diverse underwater vision tasks and foster the development of underwater camouflaged object tracking and beyond.
Our key contributions are as follows: 1) We introduce UW-COT220, the first large-scale multi-modal underwater camouflaged object tracking dataset with bounding boxes, masks, and language descriptions. 2) We propose VL-SAM2, a new VL tracker based on SAM2, leveraging language prompts and motion-aware prediction for robust underwater tracking. 3) Our VL-SAM2 outperforms state-of-the-art (SOTA) tracking methods, proving its effectiveness in handling underwater camouflage and complex scenes.
@@ 2 UW-COT220 Dataset
Our goal is to construct a large-scale underwater camouflaged object tracking dataset that involves a wide variety of categories and various real underwater scenes for evaluating and developing general underwater camouflaged object tracking methods. To achieve this, we collect underwater videos from video-sharing platforms (e.g., YouTube [^1] https://www.youtube.com/) and existing tracking datasets (e.g., WebUOT-1M [38] and VastTrack [22]), and we filter out 220 videos that contain underwater camouflaged objects. Following existing datasets [34, 22, 38], we provide a bounding box annotation for the camouflaged object in each frame, represented as $[x,y,w,h]$, where $x$ and $y$ denote the coordinates of the top-left corner of the object, and $w$ and $h$ indicate the width and height of the object. In each video, we annotate an English sentence describing the color, behavior, attributes, and surroundings of the target. Furthermore, to enhance the accuracy of object representation, we use the SAM-powered interactive semi-automatic annotation tool [17] to generate mask annotations for the camouflaged objects. Some examples and detailed statistics of the UW-COT220 dataset are provided in Fig. 1 and Tab. 1. To the best of our knowledge, UW-COT220 is the first large-scale multi-modal benchmark for underwater camouflaged object tracking, featuring a diverse set of 96 categories and various challenging underwater scenes.
3 Proposed Method
@@
In this section, we introduce vision-language (VL)-SAM2, a new vision-language tracking framework for multi-modal underwater camouflaged object tracking. VL-SAM2 (see Fig. 2) is based on the latest video foundation model, SAM2 [29], and extends its capabilities to underwater scenarios. VL-SAM2 consists of three main components: a visual branch (i.e., SAM2), a language branch (LB), and a motion-aware target prediction (MATP) module [38].
@@
The input to VL-SAM2 consists of video frames and a language description. First, we use an image encoder and a language encoder to extract features from both modalities. In the first frame, the initial bounding box is encoded by the prompt encoder to obtain the visual embedding. The language embedding is processed by a linear projection to match the dimension of the visual embedding. Then, we concatenate the visual and language embeddings as sparse embeddings and feed them into the mask decoder. In subsequent frames, we adopt a video propagation manner similar to SAM2 to obtain the prediction result for each frame. The MATP, based on Kalman filtering [18], is used to alleviate the model drift caused by similar distractors around the true target. Its core idea is to predict the current state (i.e., target box) from the previous state and then refine it by integrating current observations to determine the optimal state. Algorithm 1 presents the pseudo-code for MATP, which serves as a plug-and-play module to refine the predicted bounding boxes for more precise target boxes in subsequent frames. Inspired by [29], we further apply MATP to update the memory of SAM2.
@@
We use Hiera-L [26] as the image encoder and CLIP ViT-B/32 [23] as the language encoder. During training, we freeze the parameters of both the image encoder and the language encoder, and fine-tune the other modules. This allows us to train the proposed VL-SAM2 on an Ubuntu server with 8 A6000 GPUs. Our model was trained on the Ref-YouTube-VOS [27] dataset and evaluated in a zero-shot manner on the UW-COT220 dataset.
Algorithm 1 Inference with MATP
0:  Kalman filter $kf$, first frame, initial box, response map, candidate set $C_{t}$, maximum response set $C^{\prime}_{t}$, scores list $scores$, estimation box $b^{E}$, match result box $b^{M}$, IoU threshold $conf=0.6$, response map threshold $threshold=0.8$, IoU threshold $iou\_threshold=0.5$, match state $match\_state=False$.
0:  Target boxes ${B}$
1:  $kf$.init(first frame)
2:  $B$=[initial box]
3:  for $i=2,3,...,T$ frames do
4:     $C_{t}\leftarrow$ extract_candidates(response map, $threshold$)
5:     $C^{\prime}_{t}\leftarrow$ NMS($C_{t}$)
6:     $b^{E}\leftarrow kf$.predict()
7:     $scores\leftarrow$ compute_scores($b^{E}$, $C^{\prime}_{t}$)
8:     if iou_of(max response box, $b^{E}$) $<conf$ then
9:        $match\_state\leftarrow True$
10:     else
11:        $match\_state\leftarrow False$
12:     end if
13:     if $match\_state$ then
14:        $b^{M}\leftarrow$ argmax($scores$)
15:     else
16:        $b^{M}\leftarrow$ max response box
17:     end if
18:     $B$.append($b^{M}$)
19:     $kf$.update($b^{M}$)
20:  end for
21:  return  $B$
4 Experiments
@@
Figure 3: Comparison of SOTA trackers on the UW-COT220 dataset using AUC, Pre, cAUC, and nPre scores. Best viewed by zooming in.
Main Results. We evaluate SAM/SAM2-based tracking methods (i.e., SAM-DA [10], Tracking Anything [30], SAM2, SAMURAI-L [29], and VL-SAM2), and current SOTA VOT methods (i.e., OSTrack [31], SeqTrack [5], and ARTrack [28]) on the proposed UW-COT220. The results are shown in Fig. 3. Our observations are as follows: 1) Our VL-SAM2 achieves the best performance, surpassing the SOTA trackers, including SAMURAI-L, SeqTrack and ARTrack. 2) The top three trackers (VL-SAM2, SAMURAI-L, and SAM2) are all SAM2-based, highlighting the promising trend of using video foundation models for tracking. 3) SAM2-based methods outperform SAM-based trackers (SAM-DA and Tracking-Anything) on UW-COT220, which can be attributed to a series of improvements SAM2 introduces over SAM for video and image tasks, such as improving temporal consistency, robustness to occlusions, feature embeddings, computational efficiency, motion estimation accuracy, generalization to new domains, and integration of contextual information [25, 37].
@@
Table 2: Impact of using different points (within the initial target box) as the point prompt for SAM2. We report AUC, nPre, Pre, cAUC, and mACC scores on the UW-COT220 dataset.
  Method Point type AUC Pre nPre cAUC mACC
SAM2-B Center point 55.0 56.0 62.2 54.2 55.4
@@
Table 3: Comparison of different models for SAM2 on the UW-COT220 dataset.
  Method Size (M) Speed (FPS) AUC Pre nPre cAUC mACC
SAM2-T 38.9 47.2 53.1 53.7 60.1 52.2 53.4
SAM2-S 46.0 43.3 52.8 53.6 60.2 52.0 53.1
SAM2-B 80.8 34.8 55.0 56.0 62.2 54.2 55.4
SAM2-L 224.4 24.2 60.6 63.5 67.3 60.0 61.3
Impact of Prompt. We take SAM2 [25] as our baseline tracker to explore the impact of different ways of point prompts (i.e., center point and random point within the initial target box). From Tab. 2, we find that using the center point as a prompt yields significantly better results than using random points. This implies that for the interactive segmentation model SAM2, the quality of the input prompt plays a crucial role in determining its performance. This also motivates us to explore the use of more flexible prompts, e.g., language descriptions, in the context of underwater camouflaged object tracking.
Model Size & Speed Analysis. Tab. 3 demonstrates that larger model sizes generally lead to better performance, but the speed of the model decreases significantly. We also discovered an interesting phenomenon: when the number of model parameters is relatively small, a smaller model (e.g., SAM2-T) can even outperform a larger model (e.g., SAM2-S). We suspect this may be due to overfitting, or that relatively small models (e.g., SAM2-S) are more sensitive to the quality of the training data [25].
Impact of Language Encoder. In Tab. 4, we explore the impact of using different language encoders in the proposed VL-SAM2. T5 [24] is a large language model that models natural language processing tasks as text-to-text problems, while CLIP [23] is a contrastive language-image pre-training model. The result demonstrates that using CLIP as the language encoder outperforms T5. We hypothesize that this is due to CLIP’s text encoder being aligned with the visual encoder in the feature space over a vast amount of data, making it more suitable for the VOT task.
@@
  Method Language Encoder AUC Pre nPre cAUC mACC
VL-SAM2 T5 [24] 54.6 54.8 59.3 53.9 55.1
VL-SAM2 CLIP [23] 71.3 75.2 78.2 70.8 72.4
Table 5: Evaluation on LaSOT and WebUOT-1M datasets.
    Method LaSOT WebUOT-1M
AUC Pre nPre AUC Pre nPre
CiteTracker-256 [21] 65.9 70.2 75.7 54.6 49.3 58.4
VLTTT [13] 67.3 72.1 77.6 48.3 41.7 52.1
OSTrack [31] 69.1 75.2 78.7 56.5 52.9 61.1
MixFormerV2-B [7] 69.5 75.0 79.1 51.0 45.4 54.0
SeqTrack-B256 [5] 69.9 76.3 79.7 54.0 50.8 58.5
SAM2-L [25] 59.3 61.9 64.5 60.6 65.6 69.2
VL-SAM2 (Ours) 73.4 79.4 81.8 71.9 77.2 80.0
Generalization Analysis. To demonstrate the generalization ability of the proposed VL-SAM2, we report the results on the large-scale open-air tracking dataset, i.e., LaSOT [8], and the generic underwater object tracking benchmark, i.e., WebUOT-1M [38] (see Tab. 5). Results show that VL-SAM2 outperforms the current SOTA trackers. We believe that VL-SAM2 demonstrates strong generalization capability, ranging from challenging underwater scenarios to complex open-air environments.
Ablation Study. We conduct experiments to verify two core modules in the VL-SAM2 framework, i.e., the language branch (LB) and MATP. As shown in Fig. 4, injecting language prompts via LB significantly improves performance compared to the baseline tracker. The performance is further enhanced by using the MATP module.
Figure 4: Ablation study on the UW-COT220 dataset.